solve by elimmination method
6x - 9y = 25.5
7y - 2x = -10.5

Answer by Tobiasz(54)

ok
6x - 9y = 25.5
2x - 7y = -10.5
Lets Find y
6x - 9y = 25.5
-3(2x - 7y = -10.5)
6x - 9y = 25.5
6x + 21y = 10.5
30y = 36
Divide by 30
y+=+1.2
OK lets plug this in. to find x.
2x - 7y = -10.5
2x - 7(1.2)= -10.5
2x - 8.4 = -10.5
add 8.4 to 8.4 and -10.5
2x = -2.1
divide by 2 to both sides
x+=+-1.05
Lets check if it works!! =]


2x - 7y = 10.5
2( -1.05) - 7 ( 1.2) = -10.5
-2.1 - 8.4 = -10.5
-10.5 = -10.5
It works! Cya

Any more questions: send them to tobiasz.chris@gmail.com